You are here: Foswiki>PartnerObserving Web>ItalyQueue2010>ObsInaf20100617 (22 Jun 2010, VincenzoTest)Edit Attach

INAF LBC+LUCIFER queue observing, 2010 June16-17

Observer: V. Testa, K.Boutsia
Telescope Operator: S.Allanson
Support Astronomer: O. Kuhn
Instrument: LUCIFER1 - LBC

Summary


Data Summary

Project Objects Filter ExpTime Notes
Grazian GOODS-N 200H+K OrderSep 2.5h 1hr out of focus - NO Tell
Rigliaco Serpens Uspec, I-bessel 10 min  
Rigliaco Serpens B-bessel, V-bessel 10 min  
Rigliaco Serpens Uspec, R-bessel 20 min  

Details

02:40 - Opening the dome - starting interface

02:42 - Flats K and Ks (too low already) and J 0001-0040

02:50 - Unwrap azimuth than pointing correction

02:58 - WT10_385 0041
  • Current IE = -84.5313
  • Current CA = -61.8785
  • Corrected IE = -51.5901
  • Corrected CA = -53.8313
  • too bright flush array

03:05 - Redo correction was not perfect 0042
  • Current IE = -51.5901
  • Current CA = -53.8313
  • Corrected IE = -51.7689
  • Corrected CA = -53.2241

03:07 - Slew to BS9145 for collimation, finished at 03:12

03:13 - Slew to Telluric standard HIP62402 - acquisition
  • object almost in the mask, X + 4pix to center

03:20 - HIP62402 =0047-0048 =
  • to bright (32000) counts was DIT=30s NDIT=2

03:24 - Redo acquisition and reduce observing time 0049-0050

03:29 - HIP62402 0051-0062
  • DIT=10s, NDIT=3

03:34 - Slew to GOODS-Grazian but guide star too faint....

03:42 - Changes Guide Star.. ok!

03:46 - MOS acquisition 0067
  • Apparently no objects in the slits. Redo

04:02 - MOS acquisition again, still ref. star in the square not centered.....

04:16 - Let's try... GOODS POINT2 0071-0082
  • images look blurry

05:24 - give Y-1.5" offset back 0083-0094
  • GOODS POINT2 -given "INITIALIZE" and now spectra are sharp
  • So the instrument was out of focus, should give INITIALIZE at least at the beginning of each night
  • but unfortunately this info is not in the manual.....

06:29 - GOODS POINT2 - 3rd hr 0095-0098
  • Problems with the GCS - not guiding
  • OB aborted
  • low elevation must have caused vibration - TCS not working properly - guiding star elongated

06:52 - Take Arcs and Flats
  • Arcs: 0099-0102
  • Flats: 0103-0108

07:02 - Telluric again, this time hopefully in focus
  • Can't preset, go to a pointing star and correct pointing
  • Still not guiding, guiding star elongated, we are losing more than 30 minutes because of this!

07:30 - Collimate on Persson star, then change telluric HIP71618 at elev. = 50
  • Still cannot find guide star
  • Try to do the aquisition with no guiding - no star in the slit
  • Give up on the telluric

07:37 SWITCH TO LBC

07:45 - Switch on LBC

07:56 - 5Bias_Bino while waiting for the telescope to be ready

07:59 - WT10_343 pointing correcting and mirroradjusting
  • got another Temperature exceeded error on the blue ccd controller error
  • Mirroradjust
######################################################################
The following values are offsets which need to be applied to the
SX Global Offsets in PSF:
dX = 2.40 mm, dY = -0.77 mm, dRX = -16.09 arcsec and dRY = -50.39 arcsec
######################################################################

08:09 - Slew to M13 and dofpia
  • Blue CCD controller error, same stupid temperature error restart camera
  • RED -2759 298 -131 731 1014 166 -6 0.61 1.23
  • BLUE 23754 0 0 0 0 0 0 1.11 9.36

Housekeeping went off on its own.... strange behavior

08:18 - dofpia
  • RED 675 158 48 -39 110 -121 0 0.73 0.30
  • BLUE -788 544 526 409 1377 -64 -100 0.89 0.75
  • RED
  • BLUE blue hit limits stop and redo

08:24 - dofpia again after taking blue mirror out of limits

08:11 -
  • RED 247 87 39 100 -76 -196 0 0.65 0.19
  • BLUE -5025 1579 1536 -450 -938 421 -100 0.73 2.15
  • RED
  • BLUE -14056 0 0 0 0 0 0 0.00 5.54

Stop and clear active optics Bad pupils in the blue.

08:36 - Change target go to Serpens
  • dofpia
  • RED -243 -201 15 -268 -189 76 7 0.79 0.21
  • BLUE -3691 41 -1418 1089 878 -578 -100 0.00 1.68
  • RED -368 -8 -284 -105 6 8 -2 0.73 0.19
  • BLUE 979 -596 323 -119 -574 -669 7 0.97 0.67
  • RED 258 170 75 66 44 -30 22 0.79 0.15
  • BLUE 622 -193 -125 167 0 -377 27 1.19 0.36 * RED -284 -111 -140 8 116 57 9 0.80 0.16 * BLUE -193 447 289 -65 -137 28 26 1.22 0.20

08:47 - Serpens BV_1
  • Red trackers failed! RPC Clien/Host com.. error.
  • trun off camera, reboot cmu

09:07 - restarted and resent BV_1
  • lost focus
  • redo dofpia
  • R 38 -5 -172 -75 -219 -172 7 0.77 0.18
  • B 602 -430 -321 31 69 -355 24 1.12 0.37
  • B 354 68 108 113 -50 -137 11 1.00 0.19
  • B 108 -2 108 -124 -73 -89 5 1.14 0.13

09:16 - Serpens BV_1
  • seeing ~ 1.0

09:46 - dofpia
  • R -98 57 -89 -91 155 -24 -18 0.72 0.12
  • B 1335 35 288 4 65 -450 -29 1.02 0.61
  • R 667 -129 301 166 -142 -131 5 0.87 0.32
  • B 199 126 -296 40 -8 -144 4 1.26 0.17

09:52 - Serpens UsR_1 (2nd run)

10:23 - dofpia
  • R -44 -252 -22 -57 21 -30 -2 1.02 0.11
  • B 274 -184 -200 113 -20 -220 2 1.27 0.22

10:26 - Serpens_UsBVI_1 (2nd run) * Had to stop at 40/48 because of brightening sky

11:21 - SA110_UsB_VI

11:27 - Slew to FF.
  • BV test 2s, 1s: 4000,2000

11:32 - SkyFlat_BV
  • 6s each side

11:40 SkyFlat_UsI *1 sec- 9000, 40000

11:47 - End of night - darks with Luci

-- VincenzoTest - 17 Jun 2010
Topic revision: r6 - 22 Jun 2010, VincenzoTest
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback